Test instrument to measure magnetic field strengths are often based on the Hall Effect. In one instrument the “probe” is a 1.0-mm- thick, 6.0-mm - wide semiconductor with a charge-carrier density of 2.1x 1021 m-3 , much less than the charge-carrier density in a conductor.

Passing a 60 mA current through the probe generates a Hall voltage of 120 mV. What is the magnetic field strength?


1
Expert's answer
2021-05-24T16:52:11-0400

The Hall voltage is given by

"V_H=\frac{IB}{ned}"

where "I" is current,

"B" is magnetic field,

"n" is charge-carrier density,

"e" is charge of electron,

"d" is thickness of the "probe"


Then:


"B=\frac{V_Hned}{I}=\frac{120\cdot2.1\cdot10^{21}\cdot1.6\cdot10^{-19}\cdot10^{-3}}{60}=0.672\ T"